Brad purchased 10 shares of stock for $10 per share. He paid a $5 commission. One year later, he purchased another 10 shares at $9 per share. Again, he paid a $5 commission. In the second year, his disappoint got the best of him and he sold 10 shares ar $7 per share, paying another $5 commission. At the end of the third year, Brad liquidated his holdings, paying $5 in commissions at $8 per share What was his dollar-weighted rate of return?
Answer #1

dollar-weighted rate of return is that rate of return that will set the present values of all cash flows and terminal values equal to the value of the initial investment.

Total payments made = $10*10 + $5 + $9*10 + $5 = $200

Total inflow = $10*7 + $5 + $8*10 +$5 = $160

To equate this 200*(1+x/100) = 160

Therefore X = 0.2 or 20%